The properties of a system, during a reversible constant pressure non￾flow process at p = 1.6 bar, changed from v1 = 0.3 m3/kg, T1 = 20°C to v2 = 0.55 m3/kg, T2 = 260°C.

The specific heat of the fluid is given by

cp = 1.5+74/(T+45)kJ/kg°c

, where T is in °C.


Determine : (i) Heat added/kg ;

(ii) Work done/kg ;

(iii) Change in internal energy/kg ;

(iv) Change in enthalpy/kg


1
Expert's answer
2021-08-24T02:54:42-0400

"\upsilon_1=0.3m^{3}/kg"

T1=20°C

"\upsilon_2=0.55m^{3}/kg"

T2=260°C

p=1.6 bar

Specific heat at constant pressure,cp="(1.5+\frac{74}{T+45})kJ/kg°C"

(i)heat added/kg :

"Q=\int_{T_1}^{T_2}c_pdT"

"=\int_{20}^{260}(1.5+\frac{74}{T+45})dT"

"=|1.5T+74log_e(T+45)|_{20}^{260}"

=1.5(260-20)+74"*log_e(\frac{260+45}{20+45})"

=474.398kJ

"\therefore heat \ added/kg=474.398kJ/kg"

(ii)Work done/kg :"W=\int_{\upsilon_1}^{\upsilon_2}pdv=p(\upsilon_2-\upsilon_1)"

=1.6*105(0.55-0.3)N.m

=40*103J=40kJ

"\therefore work\ done=40kJ/kg"

(iii)change in internal energy/kg : "\u2206u=Q-W"

=474.398-40=434.398 kJ/kg

(iv)change in enthalpy/kg (for non flow process):

∆h=Q

=474.398 kJ/kg